# Hybrid Training for Enhanced Multi-task Generalization in Multi-agent Reinforcement Learning

### v1

@@ Introduction
Our approach utilizes a QMIX-style value-based MARL method, as delineated by Rashid et al. [37], integrated within the Centralized Training with Decentralized Execution (CTDE) paradigm [33] to train the high-level policy. Similar to QMIX, it tries to learn a global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ that can be decomposed into agents’ individual value functions $Q_{1}\left(\tau_{1},z_{1}\right),\ldots,Q_{n}\left(\tau_{n},z_{n}\right)$. This global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ can be trained with the squared TD loss as follows:
@@
$L_{\mathrm{TD}}\left(\theta_{v}\right)=\mathbb{E}_{\left(\boldsymbol{\tau},\boldsymbol{\tau}^{\prime}\right)\sim\mathcal{D,R}}\Bigg{[}$ $\Big{(}r+\gamma\max_{\boldsymbol{z}^{\prime}}Q_{\mathrm{tot}}\left(\boldsymbol{\tau}^{\prime},\boldsymbol{z}^{\prime};\theta_{v}^{-}\right)$
$-Q_{\mathrm{tot}}\left(\boldsymbol{\tau},\boldsymbol{z};\theta_{v}\right)\Big{)}^{2}\Bigg{]}$
@@
Following previous MARL methods [37, 20], we use $\theta_{v}$ to denote all parameters in the value networks and the action decoder, $\theta_{v}^{-}$ to denote parameters of target networks. To address potential performance bottlenecks from sub-optimal offline data in the previous skill discovery phase, which could impair cross-task performance, we incorporate the action decoder in the hybrid training stage, enhancing skill application. When estimating Q-targets, we choose the joint general skills $z^{\prime}$ by selecting each skill $z^{\prime}_{i}$ with maximal individual Q-value $Q_{i}\left(\tau^{\prime}_{i},z^{\prime}_{i}\right)$ to avoid search in the large joint skills space, as the same as [46, 37]. Finally, We adopt a mixing network to ensure that it can satisfy the individual-global-max (IGM) [51] principle which promises the action selection with individual value functions is correct.
Figure 3: Training framework during the high-level policy learning phase of HyGen. The hybrid buffer contains trajectories from online buffer $\mathcal{R}$ and offline dataset $\mathcal{D}$. The observation encoder extracts representations from local information. Meanwhile, the mixing network employs self-attention to accommodate varying input dimensions across different tasks.
@@
One challenge that remains is that we cannot directly get skills information since there are only state and joint actions recorded in both offline datasets and online replay buffers. Reusing skills calculated by the pre-trained global trajectory encoder are obtained with global information, which does not follow CTED. Consequently, we train a local observation encoder $\hat{q}(\cdot\mid\tau_{f}^{i})$, leveraging only agent $i$’s local trajectory, comprising its specific action sequences and local observations, to infer skills. The output distribution is expected to be similar to the pre-trained global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$. We calculate the KL-divergence [17] between them to update the local observation encoder as the consistent loss $L_{\mathrm{c}}$ below:
@@
$L_{\mathrm{c}}(\phi_{o})=\sum_{i=1}^{n}\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D,R}}[D_{KL}(\hat{q}(\cdot\mid\tau_{f}^{i})\left|\right|q(z_{i}|s,\mathbf{a},i))]$
@@
Task Expert Medium
BC-best UPDeT-I UPDeT-m ODIS HyGen (ours) BC-best UPDeT-I UPDeT-m ODIS HyGen (ours)
Source Tasks
3m 96.9 $\pm$ 4 71.9 $\pm$ 14 82.1 $\pm$ 10 97.4 $\pm$ 2 99.1 $\pm$ 1 65.4 $\pm$ 14 56.6 $\pm$ 14 51.2 $\pm$ 3 85.9 $\pm$ 10 91.5 $\pm$ 11
5m6m 50.4 $\pm$ 2 12.1 $\pm$ 12 17.2 $\pm$ 28 53.9 $\pm$ 5 61.2 $\pm$ 8 21.9 $\pm$ 3 5.6 $\pm$ 4 6.3 $\pm$ 4 22.7 $\pm$ 7 31.6 $\pm$ 7
9m10m 95.3 $\pm$ 1 26.6 $\pm$ 12 3.1 $\pm$ 5 80.4 $\pm$ 8 96.4 $\pm$ 3 63.8 $\pm$ 10 34.4 $\pm$ 13 28.5 $\pm$ 10 78.1 $\pm$ 3 79.2 $\pm$ 4
Unseen Tasks
4m 92.1 $\pm$ 3 28.6 $\pm$ 21 33.0 $\pm$ 27 95.3 $\pm$ 3 95.8 $\pm$ 4 48.8 $\pm$ 21 21.6 $\pm$ 17 14.1 $\pm$ 5 61.7 $\pm$ 17 91.4 $\pm$ 8
5m 87.1 $\pm$ 10 40.1 $\pm$ 25 33.6 $\pm$ 40 89.1 $\pm$ 10 99.5 $\pm$ 1 76.6 $\pm$ 14 77.4 $\pm$ 16 67.2 $\pm$ 21 85.9 $\pm$ 11 96.5 $\pm$ 6
10m 90.5 $\pm$ 3 33.9 $\pm$ 25 54.7 $\pm$ 44 93.8 $\pm$ 2 93.5 $\pm$ 5 56.2 $\pm$ 20 36.8 $\pm$ 20 32.9 $\pm$ 11 61.3 $\pm$ 11 96.4 $\pm$ 3
12m 70.8 $\pm$ 15 10.9 $\pm$ 18 17.2 $\pm$ 28 58.6 $\pm$ 11 85.2 $\pm$ 6 24.0 $\pm$ 10 4.0 $\pm$ 5 3.2 $\pm$ 3 35.9 $\pm$ 8 81.5 $\pm$ 14
7m8m 18.8 $\pm$ 3 0.8 $\pm$ 1 0.0 $\pm$ 0 25.0 $\pm$ 15 28.9 $\pm$ 12 1.6 $\pm$ 1 2.4 $\pm$ 2 0.0 $\pm$ 0 28.1 $\pm$ 22 24.5 $\pm$ 9
8m9m 15.8 $\pm$ 3 1.6 $\pm$ 1 0.0 $\pm$ 0 19.6 $\pm$ 6 25.7 $\pm$ 9 3.1 $\pm$ 3 3.1 $\pm$ 3 2.3 $\pm$ 2 4.7 $\pm$ 2 22.3 $\pm$ 10
10m11m 45.3 $\pm$ 11 0.8 $\pm$ 1 0.0 $\pm$ 0 42.2 $\pm$ 7 57.2 $\pm$ 13 19.7 $\pm$ 8 2.4 $\pm$ 1 4.0 $\pm$ 3 29.7 $\pm$ 15 47.2 $\pm$ 13
10m12m 1.0 $\pm$ 1 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 13.8 $\pm$ 4 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 5.2 $\pm$ 2
13m15m 0.0 $\pm$ 0 0.0 $\pm$ 0 0.0 $\pm$ 0 2.3 $\pm$ 2 9.5 $\pm$ 5 0.6 $\pm$ 1 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 9.3 $\pm$ 6
Table 1: Average test win rates of the final policies in the task set marine-hard with different data qualities, averaged over five random seeds. We use abbreviated names for asymmetric tasks for simplicity. For instance, ’9m10m’ refers to the SMAC map ’9m_vs_10m’. Results of BC-best represent the highest test win rates between BC-t and BC-r.
The out-of-distribution (OOD) problem refers to the challenge of dealing with situations or state-action pairs that were not encountered in the pre-collected dataset on which the agent is trained and it directly impacts the reliability and generalization of the trained models to new, unseen environments. To tackle the out-of-distribution issue, we adopt the popular conservative Q-learning (CQL) [25] method. Different from purely offline RL, in this project, the use of offline data changes according to $R_{h}$. Therefore, when using CQL loss, $R_{h}$ is used as a coefficient to control the impact of CQL on the learning process. To be concise, the total loss term in the high-level policy learning phase is presented as
$L_{\mathrm{total}}(\theta_{v},\phi_{o})=L_{\mathrm{TD}}\left(\theta_{v}\right)+\alpha\cdot L_{\mathrm{c}}(\phi_{o})+\eta\cdot R_{h}\cdot L_{\mathrm{CQL}}$
@@
In this section, we assess HyGen’s multi-task generalization capabilities, specifically focusing on zero-shot generalization across unseen tasks. Our experiments utilize custom-designed task sets from the StarCraft Multi-Agent Challenge (SMAC) [39], where we employ offline data of varied quality integrated with corresponding online interaction environments. We benchmark HyGen against purely online and offline methods across multiple source tasks, further examining its transfer capabilities in multi-scenario zero-shot transfer tasks. Experimental outcomes demonstrate that HyGen significantly outperforms purely online and offline methods.
Baselines
@@
We compare HyGen with state-of-the-art multi-task MARL methods. Given the scarcity of such methods, we include baselines developed by ODIS [60]:
  • BC-t [48, 60], a transformer-based behavior cloning method sharing the same structure as ODIS [60].
  • BC-r [48, 60], a transformer-based behavior cloning method that incorporates return-to-go information [6] in addition to the features of BC-t.
  • UPDeT-m [20, 60], a transformer-based universal MARL model using the transformer-based mixing network of ODIS [60] to facilitate simultaneous multi-task learning.
  • UPDeT-l [20, 60], a transformer-based universal MARL model that utilizes the linear decomposable network from VDN [46] for multi-task learning.
  • ODIS [60], an offline multi-task MARL method capable of discovering general skills and learning generalizable policies, thus enabling zero-shot generalization to unseen tasks.
  • QMIX [37], a prevalent online MARL baseline operating under the CTDE paradigm, lacks zero-shot generalization capabilities. However, it still can be utilized to validate the efficiency of HyGen.
@@
For our experiments in SMAC, we use the marine-hard and stalker-zealot task sets and both expert and medium offline datasets as defined and collected by ODIS [60], to ensure comparable fairness. Those task sets include three source tasks for training and multiple unseen tasks for evaluation. Agents are required to control various units, such as marines and stalkers, with the number of controllable agents and target enemies differing across tasks. Detailed descriptions of the task sets and properties of the offline datasets are available in Appendix A.
Experimental Results

### v2

@@ 4.2 Hybrid High-level Policy Learning
Our approach utilizes a QMIX-style value-based MARL method, as delineated by (rashid2020monotonic, ), integrated within the Centralized Training with Decentralized Execution (CTDE) paradigm (oliehoek2008optimal, ) to train the high-level policy. Similar to QMIX, it tries to learn a global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ that can be decomposed into agents’ individual value functions $Q_{1}\left(\tau_{1},z_{1}\right),\ldots,Q_{n}\left(\tau_{n},z_{n}\right)$. This global value function $Q_{\mathrm{tot}}(\boldsymbol{\tau},\boldsymbol{z})$ can be trained with the squared TD loss as follows:
@@
$L_{\mathrm{TD}}\left(\theta_{v}\right)=\mathbb{E}_{\left(\boldsymbol{\tau},\boldsymbol{\tau}^{\prime}\right)\sim\mathcal{D,R}}\Bigg[$ $\Big(r+\gamma\max_{\boldsymbol{z}^{\prime}}Q_{\mathrm{tot}}\left(\boldsymbol{\tau}^{\prime},\boldsymbol{z}^{\prime};\theta_{v}^{-}\right)-Q_{\mathrm{tot}}\left(\boldsymbol{\tau},\boldsymbol{z};\theta_{v}\right)\Big)^{2}\Bigg]$ (4)
@@
Following previous MARL methods (rashid2020monotonic, ; hu2021updet, ), we use $\theta_{v}$ to denote all parameters in the value networks and the action decoder, $\theta_{v}^{-}$ to denote parameters of target networks. To address potential performance bottlenecks from sub-optimal offline data in the previous skill discovery phase, which could impair cross-task performance, we incorporate the action decoder in the hybrid training stage, enhancing skill application. When estimating Q-targets, we choose the joint general skills $z^{\prime}$ by selecting each skill $z^{\prime}_{i}$ with maximal individual Q-value $Q_{i}\left(\tau^{\prime}_{i},z^{\prime}_{i}\right)$ to avoid search in the large joint skills space, as the same as (sunehag2017value, ; rashid2020monotonic, ). Finally, We adopt a mixing network to ensure that it can satisfy the individual-global-max (IGM) (wang2020qplex, ) principle which promises the action selection with individual value functions is accurate.
@@
One challenge that remains is that we cannot directly get skills information since there are only state and joint actions recorded in both offline datasets and online replay buffers. Reusing skills calculated by the pre-trained global trajectory encoder is obtained with global information, which does not follow CTDE. Consequently, we train a local observation encoder $\hat{q}(\cdot\mid\tau_{f}^{i})$, leveraging only agent $i$’s local trajectory, comprising its specific action sequences and local observations, to infer skills. The output distribution is expected to be similar to the pre-trained global trajectory encoder $q(z_{i}|s,\mathbf{a},i)$. We calculate the KL-divergence (hershey2007approximating, ) between them to update the local observation encoder as the consistent loss $L_{\mathrm{c}}$ below:
@@
$L_{\mathrm{c}}(\phi_{o})=\sum_{i=1}^{n}\mathbb{E}_{\boldsymbol{\tau}\sim\mathcal{D,R}}[D_{KL}(\hat{q}(\cdot\mid\tau_{f}^{i})\left|\right|q(z_{i}|s,\mathbf{a},i))]$ (5)
@@
The out-of-distribution (OOD) problem refers to the challenge of dealing with situations or state-action pairs that were not encountered in the pre-collected dataset on which the agent is trained and it directly impacts the reliability and generalization of the trained models to new, unseen environments. To tackle the out-of-distribution issue, we adopt the popular conservative Q-learning (CQL) (kumar2020conservative, ) method. Different from purely offline RL, in this project, the use of offline data changes according to $R_{h}$. Therefore, when using CQL loss, $R_{h}$ is used as a coefficient to control the impact of CQL on the learning process. To be concise, the total loss term in the high-level policy learning phase is presented as
$L_{\mathrm{total}}(\theta_{v},\phi_{o})=L_{\mathrm{TD}}\left(\theta_{v}\right)+\alpha\cdot L_{\mathrm{c}}(\phi_{o})+\eta\cdot R_{h}\cdot L_{\mathrm{CQL}}$ (6)
@@ 5 Experiments
In this section, we assess HyGen’s multi-task generalization capabilities, specifically focusing on zero-shot generalization across unseen tasks. Our experiments utilize custom-designed task sets from the SMAC (samvelyan2019starcraft, ), where we employ offline data of varied quality integrated with corresponding online interaction environments. We benchmark HyGen against purely online and offline methods across multiple source tasks, further examining its transfer capabilities in multi-scenario zero-shot transfer tasks. Experimental outcomes demonstrate that HyGen significantly outperforms purely online and offline methods.
5.1 Performances on Multi-task Generalization
5.1.1 Baselines
@@
We compare HyGen with state-of-the-art multi-task MARL methods. Given the scarcity of such methods, we include baselines developed by ODIS (zhang2022discovering, ):
@@
For our experiments in SMAC, we use the marine-hard and stalker-zealot task sets and both expert and medium offline datasets as defined and collected by ODIS (zhang2022discovering, ), to ensure fairness. Those task sets include three source tasks for training and multiple unseen tasks for evaluation. Agents are required to control various units, such as marines and stalkers, with the number of controllable agents and target enemies differing across tasks. Detailed descriptions of the task sets and properties of the offline datasets are available in Appendix A.
5.1.2 Experimental Results
Table 1: Average test win rates of the final policies in the task set marine-hard and stalker-zealot with different data qualities, averaged over five random seeds. The results for BC-best represent the highest test win rates between BC-t and BC-r, while those for UPDeT-best represent the highest test win rates between UPDeT-l and UPDeT-m. For ease of reference, asymmetric task names are abbreviated, with ’9m10m’ denoting the SMAC map ’9m_vs_10m’.
Task Marine-Hard-Expert Marine-Hard-Medium
BC-best UPDeT-best ODIS HiSSD HyGen (ours) BC-best UPDeT-best ODIS HiSSD HyGen (ours)
Source Tasks
3m 96.9 $\pm$ 4 82.1 $\pm$ 10 97.4 $\pm$ 2 99.5 $\pm$ 0 99.1 $\pm$ 1 65.4 $\pm$ 14 56.6 $\pm$ 14 85.9 $\pm$ 10 62.7 $\pm$ 6 91.5 $\pm$ 11
5m6m 50.4 $\pm$ 2 17.2 $\pm$ 28 53.9 $\pm$ 5 66.1 $\pm$ 7 61.2 $\pm$ 8 21.9 $\pm$ 3 26.4 $\pm$ 4 22.7 $\pm$ 7 26.4 $\pm$ 4 31.6 $\pm$ 7
9m10m 95.3 $\pm$ 1 26.6 $\pm$ 12 80.4 $\pm$ 8 95.5 $\pm$ 3 96.4 $\pm$ 3 63.8 $\pm$ 10 34.4 $\pm$ 13 78.1 $\pm$ 3 73.9 $\pm$ 2 79.2 $\pm$ 4
Unseen Tasks
4m 92.1 $\pm$ 3 33.0 $\pm$ 27 95.3 $\pm$ 3 99.2 $\pm$ 1 95.8 $\pm$ 4 48.8 $\pm$ 21 21.6 $\pm$ 17 61.7 $\pm$ 17 77.3 $\pm$ 10 91.4 $\pm$ 8
5m 87.1 $\pm$ 10 40.1 $\pm$ 25 89.1 $\pm$ 10 99.2 $\pm$ 1 99.5 $\pm$ 1 76.6 $\pm$ 14 77.4 $\pm$ 16 85.9 $\pm$ 11 88.4 $\pm$ 8 96.5 $\pm$ 6
10m 90.5 $\pm$ 3 54.7 $\pm$ 44 93.8 $\pm$ 2 98.4 $\pm$ 1 93.5 $\pm$ 5 56.2 $\pm$ 20 36.8 $\pm$ 20 61.3 $\pm$ 11 98.0 $\pm$ 0 96.4 $\pm$ 3
12m 70.8 $\pm$ 15 17.2 $\pm$ 28 58.6 $\pm$ 11 75.5 $\pm$ 20 85.2 $\pm$ 6 24.0 $\pm$ 10 4.0 $\pm$ 5 35.9 $\pm$ 8 86.4 $\pm$ 6 81.5 $\pm$ 14
7m8m 18.8 $\pm$ 3 0.8 $\pm$ 1 25.0 $\pm$ 15 35.3 $\pm$ 10 28.9 $\pm$ 12 1.6 $\pm$ 1 2.4 $\pm$ 2 28.1 $\pm$ 22 14.2 $\pm$ 10 24.5 $\pm$ 9
8m9m 15.8 $\pm$ 3 1.6 $\pm$ 1 19.6 $\pm$ 6 47.0 $\pm$ 6 25.7 $\pm$ 9 3.1 $\pm$ 3 3.1 $\pm$ 3 4.7 $\pm$ 2 15.3 $\pm$ 3 22.3 $\pm$ 10
10m11m 45.3 $\pm$ 11 0.8 $\pm$ 1 42.2 $\pm$ 7 86.3 $\pm$ 15 57.2 $\pm$ 13 19.7 $\pm$ 8 4.0 $\pm$ 3 29.7 $\pm$ 15 43.6 $\pm$ 5 47.2 $\pm$ 13
10m12m 1.0 $\pm$ 1 0.0 $\pm$ 0 1.6 $\pm$ 1 14.5 $\pm$ 9 13.8 $\pm$ 4 0.0 $\pm$ 0 0.0 $\pm$ 0 1.6 $\pm$ 1 0.6 $\pm$ 0 5.2 $\pm$ 2
13m15m 0.0 $\pm$ 0 0.0 $\pm$ 0 2.3 $\pm$ 2 1.3 $\pm$ 2 9.5 $\pm$ 5 0.6 $\pm$ 1 0.0 $\pm$ 0 1.6 $\pm$ 1 1.4 $\pm$ 2 9.3 $\pm$ 6
Task Stalker-Zealot-Expert Stalker-Zealot-Medium
Source Tasks
2s3z 93.1 $\pm$ 4 53.1 $\pm$ 39 97.7 $\pm$ 2 95.2 $\pm$ 1 97.1 $\pm$ 3 48.8 $\pm$ 9 35.0 $\pm$ 23 49.2 $\pm$ 8 32.3 $\pm$ 11.7 73.5 $\pm$ 11
2s4z 78.1 $\pm$ 8 48.4 $\pm$ 24 60.9 $\pm$ 6 79.8 $\pm$ 6 86.2 $\pm$ 10 12.5 $\pm$ 8 28.8 $\pm$ 4 32.8 $\pm$ 12 17.0 $\pm$ 2.2 51.3 $\pm$ 8
3s5z 92.5 $\pm$ 4 40.6 $\pm$ 11 87.5 $\pm$ 9 92.8 $\pm$ 5 88.9 $\pm$ 13 24.4 $\pm$ 12 25.6 $\pm$ 24 28.9 $\pm$ 6 24.4 $\pm$ 7.9 52.6 $\pm$ 13
Unseen Tasks
1s3z 45.6 $\pm$ 23 26.6 $\pm$ 25 76.6 $\pm$ 3 81.6 $\pm$ 15.2 84.1 $\pm$ 5 21.9 $\pm$ 37 33.1 $\pm$ 18 41.4 $\pm$ 18 44.2 $\pm$ 9.9 54.2 $\pm$ 8
1s4z 60.0 $\pm$ 32 37.5 $\pm$ 31 17.2 $\pm$ 10 42.0 $\pm$ 26.1 44.5 $\pm$ 9 6.2 $\pm$ 7 35.0 $\pm$ 7 50.7 $\pm$ 7 18.1 $\pm$ 11.0 67.3 $\pm$ 7
1s5z 45.6 $\pm$ 26 29.7 $\pm$ 26 2.5 $\pm$ 2 16.7 $\pm$ 12.3 47.2 $\pm$ 13 3.1 $\pm$ 2 13.1 $\pm$ 11 14.1 $\pm$ 8 2.5 $\pm$ 2.2 34.2 $\pm$ 13
2s5z 75.6 $\pm$ 11 27.3 $\pm$ 19 27.3 $\pm$ 6 79.7 $\pm$ 2.2 72.4 $\pm$ 15 14.4 $\pm$ 9 17.5 $\pm$ 9 32.0 $\pm$ 4 11.3 $\pm$ 3.7 43.7 $\pm$ 5
3s3z 80.6 $\pm$ 9 49.2 $\pm$ 25 89.1 $\pm$ 5 88.0 $\pm$ 4.5 93.3 $\pm$ 6 45.6 $\pm$ 14 24.4 $\pm$ 28 23.4 $\pm$ 9 21.9 $\pm$ 10.7 41.3 $\pm$ 7
3s4z 92.5 $\pm$ 5 59.4 $\pm$ 16 96.9 $\pm$ 2 88.1 $\pm$ 9.0 93.9 $\pm$ 5 40.0 $\pm$ 19 28.8 $\pm$ 31 50.8 $\pm$ 15 17.2 $\pm$ 4.5 71.6 $\pm$ 9
4s3z 67.5 $\pm$ 19 50.8 $\pm$ 24 64.1 $\pm$ 13 88.6 $\pm$ 4.1 74.9 $\pm$ 13 28.8 $\pm$ 26 11.2 $\pm$ 18 13.3 $\pm$ 7 31.9 $\pm$ 23.2 52.6 $\pm$ 13
4s4z 53.1 $\pm$ 18 41.4 $\pm$ 16 79.7 $\pm$ 10 73.4 $\pm$ 5.2 74.1 $\pm$ 16 20.0 $\pm$ 12 1.9 $\pm$ 2 12.5 $\pm$ 7 13.2 $\pm$ 6.5 44.2 $\pm$ 15
4s5z 40.6 $\pm$ 19 28.1 $\pm$ 17 86.7 $\pm$ 12 65.6 $\pm$ 3.7 89.9 $\pm$ 7 14.4 $\pm$ 8 5.6 $\pm$ 8 7.0 $\pm$ 4 4.5 $\pm$ 1.3 28.1 $\pm$ 13
4s6z 48.1 $\pm$ 23 10.9 $\pm$ 7 88.3 $\pm$ 8 68.4 $\pm$ 4.9 86.1 $\pm$ 12 3.8 $\pm$ 3 2.5 $\pm$ 2 1.6 $\pm$ 1 0.9 $\pm$ 0.9 23.8 $\pm$ 14